### v1

# Shortcuts to Adiabaticity via Adaptive Quantum Zeno Measurements

Abstract
We consider the quantum Zeno dynamics arising from monitoring a time-dependent projector. Starting from a stroboscopic measurement protocol, it is shown that the effective Hamiltonian for Zeno dynamics involves a nonadiabatic geometric connection that takes the form of the Kato-Avron Hamiltonian for parallel transport, stirring the evolution within the time-dependent Zeno subspace. The latter reduces to counterdiabatic driving when projective measurements are performed in the instantaneous energy eigenbasis of the quantum system. The effective Zeno Hamiltonian can also be derived in the context of continuous quantum measurements of a time-dependent observable and the non-Hermitian evolution with a complex absorbing potential varying in time. Our results thus provide a unified framework for realizing shortcuts to adiabaticity via adaptive quantum Zeno measurements.
Stirring the dynamics of a quantum system towards a target state or subspace is an ubiquitous task in quantum science and technology. Adiabatic protocols provide a natural approach to this end. However, they require slow driving, which can lead to the accumulation of uncontrollable sources of noise and errors. Shortcuts to adiabaticity provide an alternative by leading to the same target state under fast driving [10]. A wide variety of techniques fall under the umbrella of such shortcuts [40, 22, 24]. Counterdiabatic driving (CD) [13, 14, 15], also known as transitionless quantum driving [8], relies on auxiliary controls to fast-forward the adiabatic evolution. CD has been generalized to accelerate an arbitrary reference trajectory [2]. At the single-particle level, it has been experimentally demonstrated in a wide variety of systems, including nitrogen-vacancy centers [45], trapped ions [6], atomic ensembles [16], and superconducting qubits [43], among others.
The realization of CD in many-body systems is challenging as the CD term involves spatially nonlocal and multiple-body interactions [12]. This has led to the introduction of a variety of approximation schemes, including variational methods [34, 36], and expansions with controlled nonlocality [11, 38, 21]. Approximate CD schemes have given rise to a family of heuristic quantum algorithms for quantum optimization [9, 26]. Such approximate CD schemes have been implemented via digital quantum simulation [46, 27, 42]. Recent developments have shown that Trotterized dynamics can be used to directly implement the counterdiabatic time-evolution operator, circumventing the need to engineer nonlocal multiple-body CD Hamiltonian terms [25]. We also note that CD is closely related to parallel transport, which describes a driven system in which the probability amplitude in a given basis remains constant in time up to a phase factor. The generator of parallel transport has been known since early proofs of the adiabatic theorem and its generalizations [1, 7, 39].
CD techniques rely on coherent quantum control, although they can be generalized to open quantum systems with nonunitary dynamics [41, 2]. Other prominent approaches to control quantum evolution utilize quantum measurements. An instance is the celebrated quantum Zeno effect, which relies on a sequence of projective measurements, and has long been recognized as a mechanism to freeze the time evolution and the coupling to the surrounding environment [31, 28, 18]. The evolution of a system with Hamiltonian $H$, subject to frequent projective measurements associated with a projector $P$, is described by the effective Hamiltonian $H_{\rm Z}=PHP$ in the Zeno limit [20, 17]. Recently, the quantum Zeno effect was used experimentally to guide the evolution of a quantum system [23]. Such quantum Zeno dragging is achieved by continuous quantum measurements of a time-varying observable. The analogy between this approach and CD has been pointed out [30].
In this Letter, we derive the effective Hamiltonian under a sequence of time-varying projective measurements. This result generalizes the celebrated expression for quantum Zeno dynamics by including an additional geometric term, the Kato-Avron Hamiltonian that generates parallel transport. Under projective measurements in the instantaneous energy eigenbasis of the driven quantum system, the quantum Zeno Hamiltonian becomes identical to that under CD. However, the global time evolution is nonunitary and suppresses transitions and quantum coherence among the measurement eigenspaces. These results are further generalized to the case of continuous quantum measurements with a time-varying observable, as well as the case of non-Hermitian quantum evolution with a time-dependent complex absorbing potential. Thus, adaptive quantum Zeno measurements provide a versatile approach to engineer shortcuts to adiabaticity.
Stroboscopic Zeno dynamics with an adaptive projector.— Consider a quantum system with Hamiltonian $H(t)$ acting on a Hilbert space $\mathcal{H}$. We perform a sequence of projective measurements of a time-dependent projector $P(t)$ at times $t_{k}=k,\delta t$ for $k=0,\dots,N$, interspersed with unitary evolution generated by $H(t)$. Let us assume that $P(t)$ is differentiable in time, $P(t)^{2}=P(t)$ and $P(t)^{\dagger}=P(t)$, and with a constant rank in time. The stroboscopic evolution operator is
$U_{N}(T)=\prod_{k=0}^{N-1}P(t_{k+1})e^{-iH(t_{k})\delta t}P(t_{k}),$ (1)
with $T=N\delta t$ and time ordering implicit. Expanding to first order in $\delta t$, $e^{-iH(t_{k})\delta t}=\mathbb{I}-iH(t_{k})\delta t$ and $P(t_{k+1})=P(t_{k})+\dot{P}(t_{k})\delta t$. Thus, each factor becomes
$P(t_{k+1})e^{-iH(t_{k})\delta t}P(t_{k})=\bigl(P+\dot{P}\delta t\bigr)\bigl(\mathbb{I}-iH\delta t\bigr)P$
$=P-iPHP\delta t+\dot{P}P\delta t,$ (2)
where all operators are evaluated at $t_{k}$. Further note that $P\dot{P}P=0$, which follows from differentiating $P^{2}=P$, and thus $\dot{P}P=[\dot{P},P]P$. Ignoring $\mathcal{O}(\delta t^{2})$ corrections, it follows that
$P(t_{k+1})e^{-iH\delta t}P(t_{k})=P-i\left(PHP-i[\dot{P},P]\right)\delta t.$ (3)
We can now consider the corresponding quantum Zeno dynamics, by taking the limit $N\to\infty$, $\delta t\to 0$ with the total time of evolution $T$ fixed. The product formula converges to a time-ordered exponential acting within the instantaneous subspace,
$U_{\rm Z}(T)=\mathcal{T}\exp\left(-i\int_{0}^{T}H_{\mathrm{eff}}(t)dt\right)P(0),$ (4)
where the effective Hamiltonian is
$H_{\rm Z}(t)=P(t)H(t)P(t)+i[\dot{P}(t),P(t)].$ (5)
The first term is the projected physical Hamiltonian, while the second term is a geometric contribution associated with parallel transport in the moving subspace associated with $P(t)$. Indeed, this is precisely the Kato-Avron Hamiltonian introduced in proofs of the adiabatic theorem [1, 7, 39].
This result can be readily generalized to a complete orthogonal set of projectors $P_{n}(t)$} satisfying $P_{n}(t)P_{m}(t)=\delta_{nm}P_{n}(t)$, and $\sum_{n}P_{n}(t)=\mathbb{I}$. Measuring a full set of outcomes $\{P_{n}(t)\}$ at each step yields a projection-value measure (PVM), and the stroboscopic propagator becomes
$U_{N}(T)=\prod_{k=0}^{N-1}\sum_{n}P_{n}(t_{k+1})e^{-iH(t_{k})\delta t}P_{n}(t_{k}).$ (6)
By a similar analysis, one finds
$H_{\rm Z}(t)=\sum_{n}P_{n}(t)H(t)P_{n}(t)+i\sum_{n}[\dot{P}_{n}(t),P_{n}(t)].$ (7)
The effective Hamiltonian is block-diagonal in the instantaneous decomposition of the Hilbert space, with $P_{n}HP_{n}$ generating the dynamics within each sector. The connection term generates parallel transport in each subspace and transitions between sectors are thus suppressed in the Zeno limit. One can show that the leading correction for finite measurement intervals $\delta t$ induces a non-Hermitian term in the effective Hamiltonian, which becomes,
$H_{\mathrm{eff}}^{(\delta t)}(t)$ $=$ $H_{\rm Z}(t)-\frac{i\delta t}{2}\,\Gamma_{\delta t}(t)+\mathcal{O}(\delta t^{2}),$
$\Gamma_{\delta t}(t)$ $=$ $2\Big(PHQHP+P\dot{P}\dot{P}P\Big).$ (8)
The first non-Hermitian contribution, $PHQHP$, is the instantaneous generalization of the familiar static-projector leakage induced by couplings between the Zeno subspace and its complement. The second contribution, $P\dot{P}\dot{P}P$, is purely geometric and quantifies leakage caused by the motion of the projector itself; it vanishes when the subspace is time independent. Note that both contributions are positive operators. In short, leading corrections for finite $\delta t$ away from the strict Zeno limit account for leakage out of the moving Zeno subspace, without altering the form of the Hermitian effective Zeno Hamiltonian, as further detailed in the End Matter.
Relation to Counterdiabatic Driving.— An important special case arises when the monitored projectors are chosen to be the instantaneous eigenprojectors of the system Hamiltonian. In this situation, the effective Hamiltonian generated by quantum Zeno dynamics coincides with the CD Hamiltonian introduced by Demirplak and Rice [13, 14, 15] and independently by Berry [8]. Let the Hamiltonian admit the spectral decomposition $H(t)=\sum_{n}E_{n}(t)P_{n}(t)$, where $P_{n}(t)$ are orthogonal and differentiable projectors.
From the general result derived above, the effective Hamiltonian governing Zeno dynamics under frequent measurements of the projectors $P_{n}(t)$} is given by (7). Using the spectral decomposition of $H(t)$, the first term simplifies as $\sum_{n}P_{n}HP_{n}=\sum_{n}E_{n}(t)P_{n}(t)=H(t)$. As we next show, the additional contribution $A\equiv i\sum_{n}[\dot{P}_{n}(t),P_{n}(t)]$ in (7) is the CD term, also known as the adiabatic gauge potential. It is purely geometric and depends only on the motion of the spectral projectors. To appreciate this, consider the case of a nondegenerate spectrum, with unit-rank projectors $P_{n}(t)=|n(t)\rangle\langle n(t)|$. Differentiating gives $\dot{P}_{n}=|\dot{n}\rangle\langle n|+|n\rangle\langle\dot{n}|$. As a result,
$i[\dot{P}_{n},P_{n}]=i\Bigl(|\dot{n}\rangle\langle n|-|n\rangle\langle\dot{n}|\Bigr)+i\langle n|\dot{n}\rangle|n\rangle\langle n|.$ (9)
Summing over $n$ and using completeness, one finds the auxiliary CD term
$H_{\rm Z}=H(t)+i\sum_{n}\left(|\dot{n}\rangle\langle n|-\langle n|\dot{n}\rangle|n\rangle\langle n|\right).$ (10)
Equation (10) coincides with the CD Hamiltonian derived in the theory of transitionless quantum driving [13, 14, 15, 8]. This operator generates exact adiabatic transport along the instantaneous eigenstates of $H(t)$ without requiring slow driving.
Note that both the adiabatic evolution and the CD dynamics are described by (entropy-preserving) unitaries that maintain quantum coherence in the energy eigenbasis. By contrast, the Zeno implementation based on frequent nonselective measurements of the projectors $\{P_{n}(t)\}$ generates a quantum channel of the form
$\rho(t)=\sum_{n}U_{n}(t)\,P_{n}(0)\,\rho_{0}\,P_{n}(0)\,U_{n}^{\dagger}(t),$ (11)
where $U_{n}(t)$ are unitary operators acting within the corresponding subspaces. The measurement-induced pinching map $\Lambda(\rho_{0})=\sum_{n}P_{n}(0)\,\rho_{0}\,P_{n}(0)$ eliminates coherences between different sectors. Such pinching operations are purity-nonincreasing and entropy-nondecreasing, implying $\mathrm{Tr}\big[\rho(t)^{2}\big]=\sum_{n}\mathrm{Tr}\big[(P_{n}\rho_{0}P_{n})^{2}\big]\leq\mathrm{Tr}\,\rho_{0}^{2}$, and $S\big(\rho(t)\big)\geq S(\rho_{0})$, with equality holding only when the initial state $\rho_{0}$ is already block-diagonal in the decomposition defined by $\{P_{n}(0)\}$. Thus, although the quantum Zeno dynamics reproduces the same geometric generator responsible for CD in each instantaneous eigenspace, it generally realizes this transport through an incoherent, entropy-producing process. By way of example, consider an initial pure state $|\psi(0)\rangle=\sum_{n}c_{n}(0)|n(0)\rangle$. Under CD, it evolves into $|\psi(t)\rangle=\sum_{n}c_{n}(0)\exp[-i\int_{0}^{t}E_{n}(s)ds-\int_{0}^{t}\langle n(s)|\dot{n}(s)ds]|n(t)\rangle$. By contrast, under adaptive Zeno measurements, it evolves into the mixed state $\rho(t)=\sum_{n}|c_{n}(0)|^{2}P_{n}(t)$.
Shortcuts via continuous monitoring.— So far, we derived the Zeno effective Hamiltonian using a stroboscopic sequence of projective measurements. We now make contact with continuous quantum measurements [44, 29] and show how the same geometric generator emerges from the stochastic master equation (SME) in the strong-measurement limit. This provides a complementary perspective in which measurement-induced shortcuts arise from the interplay of coherent dynamics, measurement backaction, and the time dependence of the monitored subspaces.
Consider the continuous monitoring of a single Hermitian observable $X(t)=\sum_{n}x_{n}\,P_{n}(t)$, with nondegenerate eigenvalues $\{x_{n}\}$ (any distinct set suffices). Then measuring $X(t)$ is operationally equivalent to monitoring the subspaces $\mathrm{Ran}\,P_{n}(t)=\{|\psi\rangle:P_{n}(t)|\psi\rangle=|\psi\rangle\}$. The standard diffusive (homodyne-type) SME for the conditioned state $\rho_{c}(t)$ is
$d\rho_{c}$ $=$ $-i[H(t),\rho_{c}]\,dt+\kappa\,\mathcal{D}[X(t)]\,\rho_{c}\,dt$ (12)
$+\sqrt{\kappa}\,\mathcal{H}[X(t)]\,\rho_{c}\,dW(t),$
where $dW$ is a Wiener increment ($\mathbb{E}[dW]=0$, $dW^{2}=dt$), and $\kappa$ is the measurement strength. Since $X(t)$ is Hermitian, $\mathcal{D}[X]\rho=X\rho X-\tfrac{1}{2}\{X^{2},\rho\}$ and $\mathcal{H}[X]\rho=2(X\rho-\langle X\rangle\rho)$. The corresponding measurement record may be written as $dY(t)=2\sqrt{\kappa}\,\langle X(t)\rangle_{c}\,dt+dW(t)$, with $\langle\cdot\rangle_{c}=\mathrm{Tr}(\cdot\,\rho_{c})$. Averaging (12) over the measurement outcomes yields the unconditional master equation $\dot{\rho}=-i[H(t),\rho]+\kappa\,\mathcal{D}[X(t)]\,\rho$, which describes measurement-induced dephasing in the instantaneous eigenspaces of $X(t)$, i.e., the PVM $\{P_{n}(t)\}$.
To connect with the stroboscopic Zeno Hamiltonian, it is useful to introduce the unitary intertwiner $W(t)$ that transports the PVM:
$\dot{W}(t)=K(t)\,W(t),\qquad K(t)\equiv-iA=\sum_{n}[\dot{P}_{n}(t),P_{n}(t)],$ (13)
with $W(0)=\mathbb{I}$, for which one has $P_{n}(t)=W(t)P_{n}(0)W^{\dagger}(t)$. Let us define the co-moving conditioned state
$\tilde{\rho}_{c}(t)=W^{\dagger}(t)\,\rho_{c}(t)\,W(t).$ (14)
Because $W(t)$ is deterministic, the Itô transformation gives an SME of the same form, but with time-independent measurement operators. For example, for the single-observable monitoring (12), define $X_{0}\equiv X(0)=\sum_{n}x_{n}P_{n}(0)$. Then,
$d\tilde{\rho}_{c}=-i[\tilde{H}(t),\tilde{\rho}_{c}]\,dt+\kappa\,\mathcal{D}[X_{0}]\,\tilde{\rho}_{c}\,dt+\sqrt{\kappa}\,\mathcal{H}[X_{0}]\,\tilde{\rho}_{c}\,dW(t),$ (15)
with the transformed Hamiltonian
$\tilde{H}(t)$ $=$ $W^{\dagger}(t)H(t)W(t)+i\,W^{\dagger}(t)\dot{W}(t),$ (16)
$W^{\dagger}(t)\dot{W}(t)$ $=$ $W^{\dagger}(t)\!\left(\sum_{n}[\dot{P}_{n}(t),P_{n}(t)]\right)\!W(t).$
Thus, the same operator that generated the geometric term in the stroboscopic derivation appears here as the inertial (connection) term $iW^{\dagger}\dot{W}$ in the co-moving frame; see also [30].
The Zeno regime corresponds to $\kappa$ being the largest rate in the problem. In the co-moving frame (15), the dissipator $\kappa\mathcal{D}[X_{0}]$ rapidly damps the off-diagonal blocks $P_{n}(0)\tilde{\rho}_{c}P_{m}(0)$ for $n\neq m$. As $\kappa\to\infty$, the dynamics becomes confined to the block-diagonal manifold $\tilde{\rho}_{c}(t)=\sum_{n}P_{n}(0)\,\tilde{\rho}_{c}(t)\,P_{n}(0)$, and the remaining evolution within each block is generated (to leading order in $1/\kappa$) by the projected Hamiltonian $P_{n}(0)\tilde{H}(t)P_{n}(0)$. Transforming back to the laboratory frame yields an effective unitary evolution within the moving subspaces with the Hamiltonian (7), in agreement with the stroboscopic Zeno result. Further, if the monitored projectors are the nondegenerate instantaneous spectral projectors of $H(t)$, the intertwiner can be chosen to be $W(t)=\sum_{n}|n(t)\rangle\langle n(0)|$, and the effective Hamiltonian equals the CD Hamiltonian. If instead one chooses simply $X=P(t)$, a similar analysis yields the Kato-Avron Hamiltonian (5). Thus, in both the stroboscopic and the continuous-measurement formulations, ideal Zeno monitoring of the instantaneous eigenspace projectors yields transitionless driving, while for finite measurement strength $\kappa$ the SME (12) provides a systematic framework to quantify deviations from the ideal limit.
Shortcuts via time-dependent complex absorbing potentials.— Complex absorbing potentials have long been used to model detectors, e.g., in the time of arrival to a given quantum subspace [3, 4, 5, 33]. Indeed, this is one of the earliest manifestations of the quantum Zeno effect, predating the celebrated work of Misra and Sudarshan [31]. Continuous and pulse observations can thus be used to induce the quantum Zeno effect [35, 37, 19, 32]. As we next show, time-dependent complex absorbing potentials, in the Zeno limit, give rise to an effective Kato-Avron Hamiltonian, and the CD Hamiltonian as a special case. Consider a quantum system evolving under a non-Hermitian Hamiltonian
$H_{\rm nh}(t)=H(t)-i\kappa\,Q(t),$ (17)
where $H(t)=H^{\dagger}(t)$ is the physical Hamiltonian, $Q(t)$ is a projector, $Q(t)^{2}=Q(t),\qquad Q(t)=Q^{\dagger}(t)$, and $\kappa>0$ sets the strength of the absorbing potential. The complementary projector $P(t)=\mathbb{I}-Q(t)$ defines the Zeno subspace in which we aim to confine the dynamics. The Schrödinger equation reads $i\frac{d}{dt}|\psi(t)\rangle=\big[H(t)-i\kappa Q(t)\big]|\psi(t)\rangle$. For large $\kappa$, components of the wavefunction in $\mathrm{Ran}\,Q(t)$ decay rapidly, suggesting that the long-time evolution should remain confined to $\mathrm{Ran}\,P(t)$.
When the projector $P(t)$ depends on time, it is convenient to introduce a unitary operator that transports the subspaces. Define now the anti-Hermitian generator as $K(t)\equiv[\dot{P}(t),P(t)]$, and the corresponding intertwiner $W(t)$, satisfying $\dot{W}(t)=K(t)W(t)$, $W(0)=\mathbb{I}$. Using the fact that $P(t)=W(t)P(0)W^{\dagger}(t)$, $Q(t)=W(t)Q(0)W^{\dagger}(t)$, in the the co-moving frame defined by $|\tilde{\psi}(t)\rangle=W^{\dagger}(t)|\psi(t)\rangle$, the projectors become time independent. Indeed, using $\dot{W}=KW$, one finds that $|\tilde{\psi}(t)\rangle$ obeys
$i\frac{d}{dt}|\tilde{\psi}(t)\rangle=\Big(\tilde{H}(t)-i\kappa Q_{0}\Big)|\tilde{\psi}(t)\rangle,$ (18)
where $Q_{0}\equiv Q(0)$ and
$\tilde{H}(t)=W^{\dagger}(t)H(t)W(t)+iW^{\dagger}(t)\dot{W}(t).$ (19)
The second term in (19) is purely geometric and arises from the motion of the subspaces.
To perform the adiabatic elimination of the absorbing sector, we now decompose the state into components in the Zeno and absorbing subspaces as $|\tilde{\psi}(t)\rangle=|\tilde{\psi}_{P}(t)\rangle+|\tilde{\psi}_{Q}(t)\rangle$, with $|\tilde{\psi}_{P}\rangle=P_{0}|\tilde{\psi}\rangle$ and $|\tilde{\psi}_{Q}\rangle=Q_{0}|\tilde{\psi}\rangle$. Projecting Eq. (18) onto $P_{0}$ and $Q_{0}$ gives
$i\dot{\tilde{\psi}}_{P}$ $=P_{0}\tilde{H}P_{0}\,\tilde{\psi}_{P}+P_{0}\tilde{H}Q_{0}\,\tilde{\psi}_{Q},$ (20)
$i\dot{\tilde{\psi}}_{Q}$ $=Q_{0}\tilde{H}P_{0}\,\tilde{\psi}_{P}+\big(Q_{0}\tilde{H}Q_{0}-i\kappa Q_{0}\big)\tilde{\psi}_{Q}.$ (21)
For $\kappa$ large compared with the typical matrix elements of $\tilde{H}$, the component $|\tilde{\psi}_{Q}\rangle$ relaxes rapidly. To leading order, we set $\dot{\tilde{\psi}}_{Q}\simeq 0$ in Eq. (21), obtaining
$|\tilde{\psi}_{Q}\rangle\simeq\frac{1}{i\kappa}\,Q_{0}\tilde{H}P_{0}\,|\tilde{\psi}_{P}\rangle+\mathcal{O}(\kappa^{-2}).$ (22)
Substituting into Eq. (20) yields an effective Schrödinger equation $i\dot{\tilde{\psi}}_{P}=\tilde{H}_{\rm eff}(t)\,|\tilde{\psi}_{P}\rangle$, with the effective Hamiltonian
$\tilde{H}_{\rm eff}(t)=P_{0}\tilde{H}(t)P_{0}-\frac{i}{\kappa}\,P_{0}\tilde{H}(t)Q_{0}\tilde{H}(t)P_{0}+\mathcal{O}(\kappa^{-2}).$ (23)
The first term is Hermitian and generates coherent dynamics within the Zeno subspace, while the second term is anti-Hermitian and describes residual loss due to finite absorption strength.
Transforming back to the laboratory frame and keeping only the leading term in $1/\kappa$, we obtain the Kato-Avron Hamiltonian (5) as the effective Zeno Hamiltonian. Including the leading correction in $1/\kappa$ gives
$H_{\rm eff}(t)$ $=$ $PHP+i[\dot{P},P]-\frac{i}{2\kappa}\,\Gamma_{\kappa}(t)+\mathcal{O}(\kappa^{-2})$
$\Gamma_{\kappa}(t)$ $=$ $2P\big(H+i[\dot{P},P]\big)Q\big(H+i[\dot{P},P]\big)P,$ (24)
where the non-Hermitian term $\Gamma_{\kappa}(t)$ quantifies the leakage induced by finite absorption strength. As before, if $P(t)$ is chosen to be a spectral projector of $H(t)$ (or a sum of spectral projectors), the projected term satisfies $P(t)H(t)P(t)=H(t)$ within the corresponding eigenspace, and the geometric term $i[\dot{P}(t),P(t)]$ coincides with the gauge-invariant CD Hamiltonian that enforces adiabatic transport. Thus, a strong time-dependent absorbing potential provides a dynamical realization of transitionless evolution that is formally equivalent, at leading order, to both stroboscopic Zeno measurements and CD. In the static case, Schulman established a relation between the Zeno dynamics induced by stroboscopic measurements and complex absorbing potentials [35, 19, 32], which was experimentally verified [37]. Such a connection was established by analyzing the response time of a two-level system, leading to a relation between $\delta t$ and $\kappa$. As shown in the End Matter, such a relation can be generalized by comparing the generators of evolution, rather than the response time scale. Indeed, this can be done in a model-independent way for time-dependent Zeno measurements. Specifically, it can be shown that
$\Gamma_{\kappa}=\Gamma_{\delta t}+2i\Big(PHQ\dot{P}P-P\dot{P}QHP\Big).$ (25)
The additional cross terms mix dynamical and geometric couplings. These terms vanish whenever $QHP=0$ (e.g., when $P(t)$ is an exact spectral projector of $H(t)$) or when the projector is constant in time. More generally, their magnitude can be bounded as $\|\Gamma_{\rm cross}\|\leq 4\|QHP\|\,\|Q\dot{P}P\|$, showing that they are parametrically smaller than the square terms when either the Hamiltonian-induced coupling or the rate of motion of the projector is small.
In summary, we have shown that quantum control via shortcuts to adiabaticity can be induced via adaptive quantum Zeno measurements. We have shown this to be the case by considering three complementary approaches: i) pulsed measurements of a time-dependent projector, ii) continuous quantum measurements of a time-dependent observable, and iii) non-Hermitian evolution with a time-dependent complex absorbing potential. In all cases, the effective Zeno Hamiltonian yields the Kato-Avron Hamiltonian for parallel transport as a time-dependent geometric correction that steers the evolution within the varying Zeno subspace. When the Zeno measurements are associated with the spectral projectors of the instantaneous system Hamiltonian, the effective Zeno dynamics becomes identical to the counterdiabatic evolution within each sector. However, the three approaches presented are globally non-unitary and thus suppress quantum coherences between different sectors. Our results provide a unified description of shortcuts to adiabaticity enabled by quantum Zeno measurements, which can be used in combination with or as an alternative to coherent quantum control, and should be of broad interest in quantum science and technology.
Acknowledgements.— It is a pleasure to acknowledge discussions with Kasturi Ranjan Swain. This work is supported by the Luxembourg National Research Fund under Grant No. C24/MS/18940482/STAOpen.
Appendix A Finite-time-step corrections to Zeno dragging
In the ideal (mathematical) Zeno limit, $\delta t\to 0$, stroboscopic monitoring of a time-dependent projector $P(t)$ yields the effective Kato-Avron Hamiltonian within the Zeno subspace. In experiments, however, the measurement interval $\delta t$ is finite. Here we compute the leading correction in $\delta t$ by expanding all operators up to order $\delta t^{2}$. A single stroboscopic step (conditioned on staying in the Zeno subspace) from $t$ to $t+\delta t$ is described by the (unnormalized) map $|\psi(t)\rangle\ \mapsto\ |\psi(t+\delta t)\rangle=\Omega(t)\,|\psi(t)\rangle$, with $\Omega(t)\equiv P(t+\delta t)e^{-iH(t)\delta t}P(t)$, upon approximating $H(t^{\prime})\simeq H(t)$ on the interval $t,t+\delta t]$. We now expand the projector and evolution operator to second order as $P(t+\delta t)=P+\dot{P}\delta t+\frac{1}{2}\ddot{P}\delta t^{2}+\mathcal{O}(\delta t^{3})$, and $e^{-iH\delta t}=\mathbb{I}-iH\delta t-\frac{1}{2}H^{2}\delta t^{2}+\mathcal{O}(\delta t^{3})$, where all quantities without explicit time argument are evaluated at time $t$.
Multiplying the expansions and using $P^{2}=P$ gives
$\Omega(t)$ $=$ $P+\delta t\Big(\dot{P}\,P-iPHP\Big)$
$+\delta t^{2}\Big(-\tfrac{1}{2}PH^{2}P-i\dot{P}\,HP+\tfrac{1}{2}\ddot{P}\,P\Big)+\mathcal{O}(\delta t^{3}).$
Note that $\Omega(t)$ is not unitary for finite $\delta t$ and its norm loss encodes the probability of leaving the monitored subspace between measurements.
To make a separation into unitary Zeno evolution and leakage, assume the state at time $t$ lies in the Zeno subspace, $|\psi(t)\rangle=P|\psi(t)\rangle$. The survival probability after one step is
$p_{\rm surv}(t)=\frac{\langle\psi(t)|\Omega^{\dagger}(t)\Omega(t)|\psi(t)\rangle}{\langle\psi(t)|\psi(t)\rangle}.$ (27)
Using $P\dot{P}P=0$ and the identity obtained by differentiating $P\dot{P}P=0$,
$P\ddot{P}P=-2P\dot{P}\dot{P}P,$ (28)
yields the restriction of $\Omega^{\dagger}\Omega$ to the Zeno subspace up to $\mathcal{O}(\delta t^{2})$,
$P\Omega^{\dagger}\Omega P=P-\delta t^{2}\Big(PHQHP+P\dot{P}\dot{P}P\Big)+\mathcal{O}(\delta t^{3}).$ (29)
Thus, the leading leakage probability $p_{\rm leak}(t)=1-p_{\rm surv}(t)$ per step is
$p_{\rm leak}(t)=\delta t^{2}\,\frac{\langle\psi(t)|\Big(PHQHP+P\dot{P}\dot{P}P\Big)|\psi(t)\rangle}{\langle\psi(t)|\psi(t)\rangle},$
up to $\mathcal{O}(\delta t^{3})$ corrections. Equation (A) makes explicit that leakage is generically $\mathcal{O}(\delta t^{2})$ for a state initially confined to $\mathrm{Ran}P(t)$.
To connect with the Zeno Hamiltonian, it is convenient to factor the one-step map into a unitary part generated by the ideal Zeno Hamiltonian plus a small nonunitary correction. Using the effective Zeno Hamiltonian $H_{\rm Z}(t)\equiv P(t)H(t)P(t)+i[\dot{P}(t),P(t)]$, one finds that the expansion (A) can be reorganized as
$\Omega(t)$ $=$ $P(t)\left[\mathbb{I}-iH_{\rm Z}(t)\,\delta t-\frac{\delta t^{2}}{2}\,H_{\rm Z}^{2}(t)\right]P(t)$ (30)
$-\frac{(\delta t)^{2}}{2}\,\Gamma_{\delta t}(t)+\mathcal{O}(\delta t^{3}),$
where the leading nonunitary correction within the monitored dynamics is governed by the positive operator
$\Gamma_{\delta t}(t)\equiv 2\Big(PHQHP+P\dot{P}\dot{P}P\Big),$ (31)
whose expectation value controls the norm loss (A). Equivalently, the conditioned evolution within the Zeno subspace can be described by an effective non-Hermitian time-dependent Hamiltonian,
$H_{\rm eff}^{(\delta t)}(t)=H_{\rm Z}(t)-\frac{i\delta t}{2}\Gamma_{\delta t}(t)+\mathcal{O}(\delta t^{2}),$ (32)
where the Hermitian part $H_{\rm Z}(t)$ generates the ideal Zeno (parallel-transport) dynamics, while the anti-Hermitian part $\tfrac{i\delta t}{2}\Gamma_{\delta t}(t)$ quantifies leakage away from the strict Zeno limit at finite measurement intervals.
For a time-independent projector ($\dot{P}=0$), Eq. (A) reduces to the standard static-Zeno leakage $p_{\rm leak}\simeq\delta t^{2}\langle HQH\rangle$ [17]. For spectral projectors $P(t)$ of $H(t)$, the projected term $PHP$ equals the corresponding energy block, and the geometric term in $H_{\rm Z}$ coincides with the CD connection. Eq. (A) then isolates the finite $\delta t$ error due to incomplete suppression of off-block transitions.
Appendix B Relation between stroboscopic measurements and complex absorbing potentials
We have seen that the leading Zeno Hamiltonian under a complex absorbing potential can be written compactly as $H_{\rm eff}^{(\kappa)}(t)=H_{\rm Z}(t)-\frac{i}{2\kappa}\,\Gamma_{\kappa}(t)+\mathcal{O}(\kappa^{-2})$. This suggests an analogy with the finite-$\delta t$ stroboscopic result $H_{\rm eff}^{(\delta t)}(t)=H_{\rm Z}(t)-\frac{i\delta t}{2}\Gamma_{\delta t}(t)+\mathcal{O}(\delta t^{2})$, with the tentative identification $\delta t\leftrightarrow 1/\kappa$ and $\Gamma_{\delta t}\leftrightarrow\Gamma_{\kappa}$.
It is thus instructive to expand $\Gamma_{\kappa}(t)$ to compare it with $\Gamma_{\delta t}$. Using $A(t)\equiv i[\dot{P}(t),P(t)]$, one obtains
$\Gamma_{\kappa}=2P(H+A)Q(H+A)P$
$=$ $2\Big(PHQHP+PAQAP+PHQAP+PAQHP\Big).$
The second term simplifies using the projector identities $P\dot{P}P=0$ and Q=1PQ=1-P, yielding
$PAQAP=P\dot{P}\dot{P}P,$ (34)
which depends only on the motion of the subspace and has the same structure as the geometric leakage term found in the finite-$\delta t$ stroboscopic expansion.
Collecting terms, the leakage operator can therefore be written as
$\Gamma_{\kappa}=2\Big(PHQHP+P\dot{P}\dot{P}P\Big)+2i\Big(PHQ\dot{P}P-P\dot{P}QHP\Big).$ (35)
The first two contributions coincide with the operator $\Gamma_{\delta t}=2(PHQHP+P\dot{P}\,\dot{P}\,P)$ that appears in the stroboscopic Zeno expansion at finite time step $\delta t$. They represent, respectively, dynamical leakage due to couplings $QHP$ between the Zeno subspace and its complement, and geometric leakage associated with the motion of the projector. The additional cross terms mix the dynamical and geometric coupling contributions and can be bounded as discussed in the main text.
Equation (35) makes explicit the close relation between the absorber-induced Zeno dynamics and the finite-$\delta t$ stroboscopic case: in both situations, the leading correction to Zeno limit evolution is governed by a positive leakage operator built from couplings to the eliminated subspace, with small parameters $1/\kappa$ and $\delta t$, respectively. As long as the cross terms can be ignored, such leakage is formally identical in both scenarios with the identification
$\delta t=\frac{1}{\kappa}.$ (36)
This generalizes the celebrated Schulman relation [35, 37, 32], at the level of the generator of evolution, to time-dependent Zeno measurements in a model-independent way. Note that in the static case $\Gamma_{\delta t}=\Gamma_{\kappa}=2PHQHP$.

### v2

@@
 The analogy between this approach and CD has been pointed out [32].
+Adaptive Zeno measurements have also been proposed to engineer quantum heat engines [8] and for ground state preparation [31, 46].
 In this Letter, we derive the effective Hamiltonian under a sequence of time-varying projective measurements.
@@
 Ignoring $\mathcal{O}(\delta t^{2})$ corrections, it follows that
-$P(t_{k+1})e^{-iH\delta t}P(t_{k})=P-i\left(PHP-i[\dot{P},P]\right)\delta t.$ (3)
+$P(t_{k+1})e^{-iH\delta t}P(t_{k})=\left[\mathbb{I}-i\delta t\left(PHP-i[\dot{P},P]\right)\right]P.$ (3)
 We can now consider the corresponding quantum Zeno dynamics, by taking the limit $N\to\infty$, $\delta t\to 0$ with the total time of evolution $T$ fixed.
@@
 where the non-Hermitian term $\Gamma_{\kappa}(t)$ quantifies the leakage induced by finite absorption strength.
-As before, if $P(t)$ is chosen to be a spectral projector of $H(t)$ (or a sum of spectral projectors), the projected term satisfies $P(t)H(t)P(t)=H(t)$ within the corresponding eigenspace, and the geometric term $i[\dot{P}(t),P(t)]$ coincides with the gauge-invariant CD Hamiltonian that enforces adiabatic transport.
+As before, if $P(t)$ is chosen to be a spectral projector of $H(t)$ (or a sum of spectral projectors), the projected term satisfies $P(t)H(t)P(t)=H(t)$ within the corresponding eigenspace, and the geometric term $i[\dot{P}(t),P(t)]$ coincides with the gauge-invariant CD Hamiltonian that enforces adiabatic transport, as detailed in the End Matter.
 Thus, a strong time-dependent absorbing potential provides a dynamical realization of transitionless evolution that is formally equivalent, at leading order, to both stroboscopic Zeno measurements and CD.
@@
 Acknowledgements.— It is a pleasure to acknowledge discussions with Kasturi Ranjan Swain. This work is supported by the Luxembourg National Research Fund under Grant No. C24/MS/18940482/STAOpen.
-Appendix A Finite-time-step corrections to Zeno dragging
+Appendix A Operational implementation of nonselective stroboscopic PVM measurements
+In the stroboscopic construction, we repeatedly apply the completely positive map
+$\rho\mapsto\sum_{n}P_{n}(t_{k})\rho P_{n}(t_{k}),$ (26)
+corresponding to a full projection-valued measurement (PVM) $\{P_{n}(t_{k})\}$.
+This map does not require averaging over different experimental runs.
+It can be implemented on a single quantum system by coupling the system to an ancillary probe (Stinespring dilation), measuring the probe, and resetting it between steps.
+Let the probe Hilbert space admit an orthonormal pointer basis $\{|n\rangle_{A}\}$. Prepare the probe in a reference state $|0\rangle_{A}$ and apply a unitary interaction of the form
+$U_{k}=\sum_{n}P_{n}(t_{k})\otimes W_{n},\qquad W_{n}|0\rangle_{A}=|n\rangle_{A},$ (27)
+which entangles the system sectors with orthogonal probe states. Acting on an arbitrary system state $\rho$, this unitary produces
+$U_{k}\big(\rho\otimes|0\rangle\langle 0|\big)U_{k}^{\dagger}=\sum_{n,m}P_{n}(t_{k})\,\rho\,P_{m}(t_{k})\otimes|n\rangle\langle m|.$ (28)
+A projective measurement of the probe in the pointer basis $\{|n\rangle_{A}\}$ yields a definite outcome $n_{k}$ and collapses the system to $P_{n_{k}}(t_{k})\rho P_{n_{k}}(t_{k})/p_{n_{k}}$.
+The probe is then reset to $|0\rangle_{A}$, and the evolution proceeds to the next stroboscopic interval.
+However, if the measurement outcome is not used to condition subsequent control operations (nonselective protocol), the reduced state of the system after the probe measurement is obtained by tracing over the probe and results in Eq.
+(26).
+This implies discarding the classical measurement record while retaining the same physical system.
+No restart of the experiment is required and the system continues evolving within whichever sector was realized in that particular run.
+Appendix B Finite-time-step corrections to Zeno dragging
 In the ideal (mathematical) Zeno limit, $\delta t\to 0$, stroboscopic monitoring of a time-dependent projector $P(t)$ yields the effective Kato-Avron Hamiltonian within the Zeno subspace.
@@
 $p_{\rm leak}(t)=\delta t^{2}\,\frac{\langle\psi(t)|\Big(PHQHP+P\dot{P}\dot{P}P\Big)|\psi(t)\rangle}{\langle\psi(t)|\psi(t)\rangle},$
-up to $\mathcal{O}(\delta t^{3})$ corrections. Equation (A) makes explicit that leakage is generically $\mathcal{O}(\delta t^{2})$ for a state initially confined to $\mathrm{Ran}P(t)$.
+up to $\mathcal{O}(\delta t^{3})$ corrections. Equation (B) makes explicit that leakage is generically $\mathcal{O}(\delta t^{2})$ for a state initially confined to $\mathrm{Ran}P(t)$.
 To connect with the Zeno Hamiltonian, it is convenient to factor the one-step map into a unitary part generated by the ideal Zeno Hamiltonian plus a small nonunitary correction.
-Using the effective Zeno Hamiltonian $H_{\rm Z}(t)\equiv P(t)H(t)P(t)+i[\dot{P}(t),P(t)]$, one finds that the expansion (A) can be reorganized as
+Using the effective Zeno Hamiltonian $H_{\rm Z}(t)\equiv P(t)H(t)P(t)+i[\dot{P}(t),P(t)]$, one finds that the expansion (B) can be reorganized as
 $\Omega(t)$ $=$ $P(t)\left[\mathbb{I}-iH_{\rm Z}(t)\,\delta t-\frac{\delta t^{2}}{2}\,H_{\rm Z}^{2}(t)\right]P(t)$ (33)
@@
 $\Gamma_{\delta t}(t)\equiv 2\Big(PHQHP+P\dot{P}\dot{P}P\Big),$ (34)
-whose expectation value controls the norm loss (A). Equivalently, the conditioned evolution within the Zeno subspace can be described by an effective non-Hermitian time-dependent Hamiltonian,
+whose expectation value controls the norm loss (B). Equivalently, the conditioned evolution within the Zeno subspace can be described by an effective non-Hermitian time-dependent Hamiltonian,
 $H_{\rm eff}^{(\delta t)}(t)=H_{\rm Z}(t)-\frac{i\delta t}{2}\Gamma_{\delta t}(t)+\mathcal{O}(\delta t^{2}),$ (35)
@@
 For a time-independent projector ($\dot{P}=0$), Eq.
-(A) reduces to the standard static-Zeno leakage $p_{\rm leak}\simeq\delta t^{2}\langle HQH\rangle$ [17].
+(B) reduces to the standard static-Zeno leakage $p_{\rm leak}\simeq\delta t^{2}\langle HQH\rangle$ [18].
 For spectral projectors $P(t)$ of $H(t)$, the projected term $PHP$ equals the corresponding energy block, and the geometric term in $H_{\rm Z}$ coincides with the CD connection.
 Eq.
-(A) then isolates the finite $\delta t$ error due to incomplete suppression of off-block transitions.
-Appendix B Relation between stroboscopic measurements and complex absorbing potentials
+(B) then isolates the finite $\delta t$ error due to incomplete suppression of off-block transitions.
+Appendix C Multi-Sector Complex Absorbing Potentials
+Consider the non-Hermitian Hamiltonian
+$H_{\rm nh}(t)=H(t)-i\kappa\,\Lambda(t),\qquad\Lambda(t)=\sum_{n}\lambda_{n}P_{n}(t),$ (36)
+with real coefficients $\lambda_{n}$ and $\kappa>0$ large.
+If all $\lambda_{n}$ are equal, then $\Lambda(t)=\lambda\mathbb{I}$ and it induces only a global decay; therefore nontrivial sector separation requires $\lambda_{n}\neq\lambda_{m}$ for some $n\neq m$.
+In the co-moving frame $|\tilde{\psi}\rangle=W^{\dagger}|\psi\rangle$, the Schrödinger equation becomes
+$i\partial_{t}|\tilde{\psi}\rangle=\big(\tilde{H}(t)-i\kappa\Lambda_{0}\big)|\tilde{\psi}\rangle,\qquad\Lambda_{0}=\sum_{n}\lambda_{n}P_{n}(0),$ (37)
+with $\tilde{H}(t)=W^{\dagger}HW+iW^{\dagger}\dot{W}$. Decomposing $|\tilde{\psi}\rangle=\sum_{n}P_{n}(0)|\tilde{\psi}\rangle\equiv\sum_{n}|\tilde{\psi}_{n}\rangle$ yields
+$i\partial_{t}|\tilde{\psi}_{n}\rangle=\sum_{m}\tilde{H}_{nm}(t)|\tilde{\psi}_{m}\rangle-i\kappa\lambda_{n}|\tilde{\psi}_{n}\rangle,$ (38)
+where $\tilde{H}_{nm}=P_{n}(0)\tilde{H}P_{m}(0)$.
+For $n\neq m$, transitions are suppressed when $\kappa|\lambda_{n}-\lambda_{m}|\gg\|\tilde{H}_{nm}(t)\|.$ Under this condition, adiabatic elimination of off-diagonal sectors gives a leading-order effective generator
+$H_{\rm eff}(t)=H_{Z}(t)-i\kappa\sum_{n}\lambda_{n}P_{n}(t)+\mathcal{O}(\kappa^{-1}),$ (39)
+which is block diagonal in the monitored decomposition.
+A special case is the two-sector decomposition defined by a single projector $P(t)$ and its complement $Q(t)=\mathbb{I}-P(t)$.
+Choosing $\lambda_{P}=0$, $\lambda_{Q}=1$, one recovers $H_{\rm nh}(t)=H(t)-i\kappa Q(t)$, discussed in the main text.
+In this case, the effective Zeno Hamiltonian becomes $H_{\rm eff}(t)=P(t)H(t)P(t)+i[\dot{P}(t),P(t)]+\mathcal{O}(\kappa^{-1})$, demonstrating that the multi-sector construction reduces to the standard Kato–Avron Hamiltonian for a protected subspace.
+Appendix D Relation between stroboscopic measurements and complex absorbing potentials
 We have seen that the leading Zeno Hamiltonian under a complex absorbing potential can be written compactly as $H_{\rm eff}^{(\kappa)}(t)=H_{\rm Z}(t)-\frac{i}{2\kappa}\,\Gamma_{\kappa}(t)+\mathcal{O}(\kappa^{-2})$.